solve triangle XYZ where angle X=40, x=12 and y=6
Found 2 solutions by stanbon, tommyt3rd:
Answer by stanbon(75887)

solve triangle XYZ where angle X=40, x=12 and y=6
--------
sin(Y)/y = sin(X)/x
-----
sin(Y) = 6[sin(40)/12] = (1/2)sin(40) = 0.3214
----
Y = arcsin(0.3214) = 18.75 degrees
---
Z = 180-(18.75+40) = 121.25 degrees
==================
z/sin(Z) = x/sin(X)
z = sin(Z)[x/sin(X)]
z = 0.8549(12/0.6428) = 15.96
====================================
Cheers,
Stan H.

Answer by tommyt3rd(5050)

Law of Sines: sinY=(6/12)sin40
sinY=0.3214
Y=arcsine(0.3214)
Y=18.747

Since all angles must add to 180 degrees:
Z=180-40-18.747=121.25
now we find z...
z=12(sin121.25/sin40)=15.96